OpenStudy (anonymous):

What is the eccentricity of the ellipse given by the equation = x^2/49 + y^2/81 =1?

OpenStudy (dumbcow):

e = c/a a = 9 b=7 a^2 = b^2 + c^2 81 = 49 + c^2 c^2 = 32 c = 4sqrt2 \[e = \frac{4\sqrt{2}}{9}\]

OpenStudy (anonymous):

0.63 0.81 1.59 1.24

OpenStudy (anonymous):

^^^^ choises

OpenStudy (dumbcow):

ok well which one matches with my answer...use a calculator

OpenStudy (dumbcow):

here is a ref sheet for these types of equations/graphs, hopefully will help you

OpenStudy (anonymous):

you do ralize you couldve just said the right answer?

OpenStudy (dumbcow):

yes but you should learn how to use your calculator and reason out the answer on your own...i'm here to give help and giudance

OpenStudy (anonymous):

dont have a calculator